# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@ 3.5. Hybrid Approaches
While most of the models presented so far have been developed for use in a supervised setting, in recent years, various novel approaches have emerged that extend the FSL paradigm to unsupervised learning. At the same time, the close relationship of FSL with learning paradigms such as self-supervised learning or multi-task learning has led to the development of new hybrid methods that combine the best of both worlds.
3.5.1. Semi-Supervised FSL
@@
One of the first works to incorporate unlabeled data within each task’s support set was in (Ren et al., 2018), where the authors propose extending prototypical networks for semi-supervised FSL. Specifically, soft k-means clustering is used, where the initial clusters are the prototypes derived by using the labeled samples, and the refined prototypes are obtained by incorporating the clustered unlabeled samples. The framework is also extended to include unlabeled samples belonging to distractor classes by introducing an additional cluster to deal with distractor samples and proposing a soft masking mechanism in the case of multiple distractor classes. More recently, semi-supervised FSL has also been applied to tabular data (Nam et al., 2023). The proposed method, STUNT, leverages a large unlabeled dataset to create diverse tasks with pseudo-labels using k-means clustering and feature perturbations. Inference is then performed on a small labeled dataset by a Prototypical Network trained on the generated tasks.
3.5.2. Self-Supervised FSL
Although meta-learning and self-supervised learning constitute different learning paradigms and are generally used for different problems, they share a common underlying goal: to extract knowledge from upstream tasks and apply it to learning downstream tasks more efficiently. Consequently, various approaches leverage that close relationship (Ni et al., 2021) between these two learning paradigms to improve FSL, particularly in the unsupervised setting. In general, many of these techniques use an unlabeled dataset to formulate few-shot tasks that are used to train a model that performs FSL in the labeled dataset. For instance, in (Khodadadeh et al., 2019), one-shot tasks are generated by assuming that each selected sample belongs to a different class, and the corresponding query set is created by applying basic image DA techniques. Figure 9 presents a high-level overview of the model’s architecture. Additionally, in (Jang et al., 2023), each unlabeled sample is considered a query, and the top-k similar samples are selected to be in the support set using a contrastive loss. On the other hand, in (Lee et al., 2023b), task generation from unlabeled samples has been approached through set representation learning. Specifically, a contrastive loss that considers instance-level, set-level, and cross-similarity between positive pairs and instances is used during meta-training. At test time, a representation is learned for each set, which is then used as initial weights for a linear classifier that is fine-tuned on each task’s support set. Lastly, contrastive learning has been employed in the context of few-shot classification of partial views of images(Jelley et al., 2023). In that case, the partial image views are used as samples within each task, and each image constitutes a different class. Finally, a generalized version of Prototypical Networks that uses a product of experts and leverages contrastive learning is used for classification.
Figure 8. UMTRA architecture(Khodadadeh et al., 2019).
Figure 9. KML architecture(Abdollahzadeh et al., 2021).
3.5.3. Unsupervised FSL
@@
Most approaches in that setting combine unsupervised learning and meta-learning to enable rapid learning within different downstream tasks with a few labeled samples. These methods generally utilize information extracted from unlabeled data and learn how to leverage the underlying shared structure across different tasks. A common approach to dealing with unsupervised FSL is creating synthetic tasks with unlabeled data to enable meta-learning on them. CACTUS(Hsu et al., 2019), one of the first methods in that direction, performs multiple executions of k-means in the feature representation space to obtain partitions that can be used as pseudo-classes, and uses these partitions to construct tasks used in training supervised meta-learning models. In (Khodadadeh et al., 2021), a generic, domain-independent approach based on generative models is proposed for task generation. These models are trained to model the unlabeled dataset’s latent space, and same-class samples are created by sampling in the proximity of a point in the latent space. Finally, apart from task generation, other works have also focused on directly learning the multi-modality within each randomly sampled task. In (Kong et al., 2021), a top-down generative model that learns unknown labels through categorical latent variables is proposed, that consists of an Energy-Based Model (EBL) that couples latent representations with class labels. Interestingly, the label can be inferred from the latent representations, allowing the model to learn the structure of different classes for different tasks, even without access to class labels.
3.5.4. Multi-Task FSL
@@
Another combination of approaches in FSL includes that of multi-task learning (MTL) and meta-learning. In general, meta-learning can be viewed as a generalized version of MTL, aiming to generalize to previously unknown tasks. However, MTL can also provide valuable elements and techniques that meta-learning can benefit from. For instance, in (Abdollahzadeh et al., 2021), the idea of transference from MTL is extended to quantify the transference of knowledge from training tasks to test tasks in meta-learning. Additionally, a modulation mechanism (KML) based on hard parameter sharing that extends that of Multimodal MAML(Vuorio et al., 2019) to mitigate negative transfer between tasks is proposed. A high-level overview of KML’s architecture can be seen in Figure 9. Moreover, in (Ye et al., 2021), MAML’s bi-level optimization formulation is extended by introducing multiple objective functions at the upper level, and the Multiple Gradient Descent Algorithm (MGDA) is used to tackle the existence of multiple objectives. Finally, in (Yu et al., 2022), meta-learning is cast as a multi-objective optimization problem, where each task’s loss is handled as a different objective. Notably, this method can help mitigate the compromising phenomenon, where the solution focuses on minimizing the losses of only a small subset of tasks and converges faster than MAML.
4. Few-Shot Learning Applications
@@ 4.1. Computer Vision
Traditionally, most of the research pertinent to FSL has focused on computer vision problems and specifically image classification(Snell et al., 2017; Finn et al., 2017), primarily due to the existence of appropriate FSL datasets for training and evaluation. However, in recent years, novel FSL approaches have been developed aiming to target different compute vision problems, including image segmentation(Gkanatsios et al., 2023), object detection (Demirel et al., 2023) and 3D reconstruction(Kwak et al., 2023). Finally, the usage of FSL has also been extended to video inputs(Yan et al., 2023).
4.2. Reinforcement Learning
@@
One of the first applications for many of the aforementioned meta-learning techniques has been meta-RL, which involves learning to quickly adapt across various similar environments by finding an optimal policy across tasks, given only a few episodes in each one of them. This is a particularly useful skill for agents acting in dynamic environments, largely resembling human-like learning, and has been used in the context of imitation learning(Rakelly et al., 2019), visual navigation(Ni et al., 2023), multi-armed bandits(Sung et al., 2017), robot arm manipulation(Wang et al., 2023b) and control(Ni et al., 2023). Novel applications of meta-RL also include online meta-RL(Khattar et al., 2022), black-box optimization(Ma et al., 2024), long-term memory from visual observations(Beck et al., 2024) and implicit language learning(Liu et al., 2023).
4.3. Natural Language Processing
@@
Another field of research where FSL has drawn increasing attention in the last few years is that of NLP, with some standard applications including machine translation(Garcia et al., 2023), and text(Han et al., 2023) and sentiment(Yu et al., 2018) classification. More recently, FSL has also been extended to novel applications and tasks such as prompt tuning (Jiang et al., 2023b), visual question answering (Najdenkoska et al., 2023), and dense retrieval (Dai et al., 2022). Finally, based on LLMs’ ability to handle multiple downstream tasks when given only a few examples in the form of prompts (see In-context Learning in Large Language Models in Section 5), various similar FSL techniques have also been proposed for applications such as natural language inference including multiple downstream tasks like causal reasoning, textual entailment, word sense disambiguation, and question answering (Murty et al., 2021).
4.4. Audio

### v3

@@ 4 Few-Shot Learning Beyond Supervised Learning Settings
Something that can be easily inferred from the models presented so far is that almost all of these methods have been developed for use in a supervised setting. However, this greatly limits the applications in which FSL can be used since many ML applications rely on unlabeled data. As a result, various novel approaches have emerged in recent years that aim to address this limitation by extending the FSL paradigm to different types of problems. At the same time, the close relationship of FSL models with techniques developed within other learning paradigms, such as self-supervised learning or multi-task learning, has led to the bridging of these approaches and the development of new hybrid methods that combine the best of both worlds.
4.1 Semi-Supervised FSL
Figure 9: Overview of Semi-Supervised FSL, where each task’s support set includes both labeled and unlabeled samples
The FSL problem setting presented in Section 2.1 assumes that all samples in each task’s support set are labeled. Specifically, for any task $T$ in the meta-train set and meta-test set the standard FSL setting assumes that a task’s $T$ support set $D_{T}^{S}$ consists of labeled samples, i.e., $D_{T}^{S}=\{(x_{i},y_{i})\}_{i=0}^{N_{T_{L}}}$, where $N_{T_{L}}$ is the total number of labeled samples in the support set. In contrast, semi-supervised FSL addresses a more generalized setting that is more eminent under real-world conditions, where the support set includes both labeled samples $D_{l}=\{(x_{l,i},y_{l,i})\}_{i=1}^{N_{l}}$ and unlabeled samples $D_{u}=\{x_{u,i}\}_{i=1}^{N_{u}}$. In this context, the support set can be expressed as $D_{T}^{S}=D_{l}\cup D_{u}$. It is also worth noting that $N_{l}$, which is the number of labeled samples in the support set, is typically small, while $N_{u}$, which is the number of unlabeled samples in the support set, is significantly larger. Figure 9 provides an overview of the above problem formulation. Overall, the objective of semi-supervised FSL is to improve performance for downstream FSL tasks by leveraging both the labeled and unlabeled data of each task.
@@ 4.1 Semi-Supervised FSL
One of the first works to incorporate unlabeled data within each task’s support set was in [174], where the authors propose extending prototypical networks for semi-supervised FSL. Specifically, soft k-means clustering is used, where the initial clusters are the prototypes derived by using the labeled samples, and the refined prototypes are obtained by incorporating the clustered unlabeled samples. The framework is also extended to include unlabeled samples belonging to distractor classes by introducing an additional cluster to deal with distractor samples and proposing a soft masking mechanism in the case of multiple distractor classes.
In general, similar to supervised learning, directly adapting approaches commonly used in standard semi-supervised learning formulations to the FSL setting would yield inferior results due to the limited availability of labeled training samples within each task. In that setting, a common approach to effectively incorporate the available unlabeled data in each task is by propagating labels from the available labeled support set data to the unlabeled data. Specifically, some common approaches have focused on embedding both labeled and unlabeled samples into a graph and leveraging the manifold structure of the data to assign pseudo-labels to the unlabeled samples [118, 135], thus increasing the number of labeled samples available for adaptation within each task.
However, predicting pseudo-labels for unlabeled samples can be challenging, especially when only a limited amount of labeled data is available. To tackle this issue, the authors in [223] proposed reversing the problem of pseudo-labeling prediction to predicting the negative pseudo-labels for each unlabeled sample based on the class that is less probable to belong to. The method is applied iteratively to all unlabeled samples, and the actual positive pseudo-labels are eventually retrieved by eliminating all assigned negative pseudo-labels for each sample.
4.2 Unsupervised FSL
Figure 10: In the Unsupervised FSL setting, meta-train set tasks are generated using pseudo-labels assigned to the unlabeled samples available during meta-training
A major principle underlying all FSL approaches, regardless of their specific formulation, is their ability to leverage prior knowledge gained from different tasks or data sources to adapt to novel tasks quickly using only a small number of training samples. In this context, unsupervised FSL can be seen as a natural extension of supervised and semi-supervised FSL, where a large amount of data is available for meta-training, yet none of it is labeled. More specifically, in unsupervised FSL we consider that the $meta-train$ set contains only unlabeled samples, $D_{meta-train}=\{x_{i}\}_{i=1}^{N_{total}}$, where $N_{total}$ is the total number of unlabeled samples. Since the classes of the tasks in the $meta-test$ set are unknown during meta-training, it is impossible to propagate pseudo-labels from the labeled samples in the $meta-test$ set to the unlabeled samples. Consequently, techniques from semi-supervised learning cannot be successfully applied in this setting, necessitating a different approach to extracting prior knowledge from the unlabeled samples. Figure 10 provides a high-level overview of the task-generating procedure during meta-training in the Unsupervised FSL problem formulation.
@@ 4.2 Unsupervised FSL
For this specified setting, most approaches combine unsupervised learning and meta-learning to enable rapid learning within different downstream tasks with a few labeled samples. These methods generally utilize information extracted from the unlabeled data and learn how to leverage the underlying shared structure across different tasks. A common approach to dealing with unsupervised FSL is creating synthetic tasks with unlabeled data to enable meta-learning on them. CACTUS [80], one of the first methods in that direction, performs multiple executions of k-means in the feature representation space to obtain partitions that can be used as pseudo-classes, and uses these partitions to construct tasks used in training supervised meta-learning models. In [104], a generic, domain-independent approach based on generative models is proposed for task generation. These models are trained to model the unlabeled dataset’s latent space, and same-class samples are created by sampling in the proximity of a point in the latent space. Similar approaches for unsupervised FSL have also been applied to tabular data [150]. The proposed method, STUNT, leverages a large unlabeled dataset to create diverse tasks with pseudo-labels using k-means clustering and feature perturbations. Inference is then performed on a small labeled dataset by a Prototypical Network trained on the generated tasks.
Finally, apart from task generation, other works have also focused on directly learning the multi-modality within each randomly sampled task. In [111], a top-down generative model that learns unknown labels through categorical latent variables is proposed, that consists of an Energy-Based Model (EBL) that couples latent representations with class labels. Interestingly, the label can be inferred from the latent representations, allowing the model to learn the structure of different classes for different tasks, even without access to class labels.
4.3 Self-Supervised FSL
Although meta-learning and self-supervised learning constitute different learning paradigms and are generally used for different problems, they share a common underlying goal: to extract knowledge from upstream tasks and apply it to learning downstream tasks more efficiently. Additionally, similar to the methods for unsupervised FSL, self-supervised learning utilizes unlabeled data during training and uses them to create pretext tasks to train the models and improve their performance in downstream tasks with limited labeled samples. Consequently, various approaches leverage that close relationship [153] between these learning paradigms to improve FSL, particularly in the unsupervised setting.
In general, many of these techniques use an unlabeled dataset to formulate few-shot tasks that are used to train a model that performs FSL in the labeled dataset. For instance, in [103], one-shot tasks are generated by assuming that each selected sample belongs to a different class, and the corresponding query set is created by applying basic image DA techniques. Figure 12 presents a high-level overview of the model’s architecture. Additionally, in [87], each unlabeled sample is considered a query, and the top-k similar samples are selected to be in the support set using a contrastive loss. On the other hand, in [119], task generation from unlabeled samples has been approached through set representation learning. Specifically, a contrastive loss that considers instance-level, set-level, and cross-similarity between positive pairs and instances is used during meta-training. At test time, a representation is learned for each set, which is then used as initial weights for a linear classifier that is fine-tuned on each task’s support set.
Lastly, contrastive learning has been employed in the context of few-shot classification of partial views of images [88]. In that case, the partial image views are used as samples within each task, and each image constitutes a different class. Finally, a generalized version of Prototypical Networks that uses a product of experts and leverages contrastive learning is used for classification.
Figure 11: In the UMTRA architecture [103], one-shot tasks are generated by matching support set samples with their augmented counterparts in the query set
Figure 12: In the KML architecture [2], the model parameters are modulated based on the concept of transference from multi-task learning
4.4 Multi-Task FSL
Another combination of approaches in FSL includes that of multi-task learning (MTL) and meta-learning. In general, meta-learning can be seen as a more generalized version of MTL, aiming to adapt to previously unknown tasks. In MTL, a specific set of tasks is defined during training, and the model is trained to generalize to these tasks simultaneously. In contrast, in standard FSL and meta-learning settings, the tasks seen during meta-training and meta-testing differ, and the number of tasks in meta-testing may also vary.
Yet, despite these differences, MTL can provide valuable elements and techniques that meta-learning can benefit from. For instance, in [2], the idea of transference from MTL is extended to quantify the transference of knowledge from training tasks to test tasks in meta-learning. Additionally, a modulation mechanism (KML) based on hard parameter sharing that extends that of Multimodal MAML [207] to mitigate negative transfer between tasks is proposed. A high-level overview of KML’s architecture can be seen in Figure 12.
@@ 4.4 Multi-Task FSL
Moreover, in [238], MAML’s bi-level optimization formulation is extended by introducing multiple objective functions at the upper level, and the Multiple Gradient Descent Algorithm (MGDA) is used to tackle the existence of multiple objectives. Finally, in [244], meta-learning is cast as a multi-objective optimization problem, where each task’s loss is handled as a different objective. Notably, this method can help mitigate the compromising phenomenon, where the solution focuses on minimizing the losses of only a small subset of tasks and converges faster than MAML.
4.5 Cross-Domain FSL
Figure 13: Overview of Cross-Domain FSL, where the meta-train set tasks and meta-test set tasks are generated from different task-generating distributions
Cross-Domain FSL (CD-FSL), which combines FSL’s knowledge transfer across tasks with different classes with Domain Adaptation’s knowledge transfer across different data distributions, has emerged as a new learning paradigm in recent years. More specifically, CD-FSL aims to leverage knowledge extracted from the base tasks to adapt to new tasks that have both novel classes and also follow a different distribution [229].
In the problem formulation of FSL in Section 2.1, an assumption was made that all tasks in both meta-train and meta-test sets are derived from the same task distribution $q(T)$. However, in real-world applications, it is possible for the tasks in the meta-test set to originate from a different task distribution compared to those in the meta-train set. Consequently, in CD-FSL, the tasks that belong to the meta-train set $D_{meta-train}=\{D_{i}\}_{i=1}^{N}$, where $D_{i}$ is the dataset corresponding to a task $T_{i}$ are drawn from a task distribution $q_{meta-train}(T)$, i.e., $T_{i}\sim q_{meta-train}(T)$, while the tasks that belong to the meta-test set $D_{meta-test}=\{D_{j}\}_{j=1}^{M}$ are drawn from a task distribution $q_{meta-test}(T)$, i.e., $T_{j}\sim q_{meta-test}(T)$, and $q_{meta-train}\neq q_{meta-test}$. Figure 13 illustrates more clearly how tasks in $D_{meta-train}$ and $D_{meta-test}$ originate from different distributions. Although more challenging, this setting is particularly relevant since, in practice, it is typical for the base and target datasets of FSL to originate from different domains [256]. For instance, a model might be trained on a large dataset of natural images, such as ImageNet, and be adapted and evaluated on a smaller dataset that is fundamentally different, such as medical or satellite images.
A common challenge in CD-FSL is that representations on the target dataset are often biased due to the data volume discrepancy, making it difficult to obtain robust models across different domains. To address this issue, recent approaches have focused on mitigating cross-domain bias during both the development and evaluation stages. At the development level, CD-FSL includes models that are explicitly trained to handle few-shot domain adaptation/generalization, with methods such as finetuning a model’s projection and prediction heads [87], retraining the feature extractor using an inverted Firth Bias regularizer [209], and using knowledge distillation [256]. The evaluation level includes few-shot cross-domain adaptation of models, even if they were not explicitly trained for this task, which is useful in a wide variety of applications such as visual question answering, temporal point processes, neural architecture search, and drug discovery [179].
Overall, robustness against domain shifts can be a potential step towards achieving Broad AI [179]. At the same time, cross-domain evaluation in meta-learning could help measure overfitting at the meta-level by attributing generalization to learning or memorization at the task level.
4.6 Few-Shot Federated Learning
Figure 14: Overview of Few-Shot Federated Learning, where the standard FSL framework is extended to consider each client as a separate task
In recent years, decentralized types of learning like Federated Learning [9] (FL) have been extensively researched to leverage scattered data across multiple machines while preserving privacy. In the standard FL setup, there are multiple clients, each having its own private dataset, denoted as $D_{i}$ for the $i$-th client. Under this setup, the objective is to learn a global model that performs well across all clients’ datasets. To achieve this, various approaches have been developed that aim to aggregate each client’s model weights to create a global model that generalizes to all clients’ data simultaneously. This way, FL reduces communication costs and enhances data privacy by sharing only model parameters between servers and clients.
Given that there is sufficient data within each client’s dataset, it has been demonstrated that these methods can yield strong results. However, FL’s performance significantly declines in scenarios where data is limited for each client. To this end, Few-Shot Federated Learning (FS-FL) has been explored as a potential way to address these issues arising in federated applications with limited data. In FS-FL, each client is treated as a separate task with its own associated dataset. Consequently, the problem can be recast as an FSL problem that follows the same formulation as the one outlined in Section 2.1, with the additional consideration that there can be no data exchange between different tasks. Under this formulation, the FL’s goal of learning a model that performs well in all clients’ data is also recast as learning a model that can rapidly adapt to each client’s dataset with minimal training samples. Figure 14 provides an overview of the FS-FL’s problem formulation. Based on this recasting of FS-FL as an FSL problem, several approaches combining FL with meta-learning have been proposed, such as adjusting popular meta-learning algorithms for FL [23], recasting FedAvg as a meta-learning algorithm for personalized FL [95] and separately executing meta-learning’s optimization loops on the server and clients [50].
In the last few years, a plethora of different FS-FL methods have been developed for applications such as facial expression recognition [184], load forecasting [197], 3D MRI medical image classification [76], fault diagnosis in industrial settings [26], and NLP [42]. A common theme across all of these applications is the sensitive nature of the involved data that necessitates privacy preservation, as well as the fact that this data is typically scarce, heterogeneous, and dispersed in different machines. Consequently, FS-FL is an ideal solution for these types of problems.
4.7 Continual FSL
Figure 15: In the Continual FSL framework, each task appears sequentially in a session-by-session manner
Following the recent trend of bridging the gap between FSL and related fields of study, FSL has also been applied in the context of Continual Learning [37]. Continual Learning refers to the ability of a model to learn from an infinite stream of data, acquiring new knowledge along the way while also preserving previously gained information. In general, training in this setting is performed in a session-by-session manner, and one of the main challenges is overcoming catastrophic forgetting where the acquisition of new knowledge during subsequent sessions results in performance degradation on the knowledge accumulated during previous sessions. In the case of Continual FSL, each session consists of a single or a batch of different N-way-K-shot tasks with limited training samples. This scarcity of training data introduces an additional layer of difficulty since typical Continual Learning methods are susceptible to overfitting the new data. Figure 15 illustrates the session-by-session learning procedure that is typical in Continual FSL.
One of the first attempts to address Continual FSL recasts meta-learning in the online learning setting by developing a new algorithm called follow the meta-leader [56] (FTML). FTML is the meta-learning version of the widely used follow-the-leader algorithm and extends MAML [54] in the online learning setting. However, catastrophic forgetting is not explicitly addressed since all previous data is stored and is available in each session. Overfitting is dealt with by leveraging meta-learning, thus obviating the need to take many gradient steps to refine the model parameters within each task. Other notable approaches in Continual FSL include introducing a mixture of across-tasks shared parameters through a Dirichlet process mixture model to create a meta-learning model capable of incrementally learning heterogeneous tasks [91] and using Bayesian meta-learning with different meta-knowledge components encapsulating across-tasks knowledge to enable continual learning of sequentially available heterogeneous tasks [226].
4.8 Class-Incremental FSL
Figure 16: Overview of Class-Incremental FSL where in each session additional classes are added, with only limited support set samples available, and the model is tasked with predicting from all classes seen up to that point
Much of the recent research revolving around Continual FSL has mainly focused on Class Incremental FSL [250] (CIFSL) where in each session, a new set of classes is introduced, and the learning model should be able to infer from all classes encountered up to that point. This type of problem differs from the Continual FSL paradigm in that the first session (base session) typically consists of a task with a sufficiently large number of training samples. In contrast, the subsequent tasks in the incremental sessions are structured as N-way-K-shot tasks with only a limited number of samples. Figure 16 illustrates the incremental nature of the Class-Incremental FSL’s problem formulation.
In [189], a contrastive supervised loss is used to ensure adequate separation of the base classes and facilitate generalization to novel classes during the incremental sessions. Another common approach in class incremental learning is to leverage knowledge distillation to adapt to new sessions. However, standard knowledge distillation methods fail in the FSL setting as they tend to overfit novel classes with only a few samples. To mitigate this issue, a combination of two teachers has been proposed to generate the student network: one derived from the base session capturing base knowledge and one from the previous session capturing newly acquired knowledge [252]. Finally, various recently introduced methods for CIFSL draw inspiration from techniques and concepts in other areas of machine learning, such as recursive analytical learning [259], neural collapse [235], and the Lottery Ticket Hypothesis [99].
5 Few-Shot Learning Applications
@@ 5.1 Computer Vision
Traditionally, most of the research pertinent to FSL has focused on computer vision problems and specifically image classification [187, 54], primarily due to the existence of appropriate FSL datasets for training and evaluation. However, in recent years, novel FSL approaches have been developed aiming to target different compute vision problems, including image segmentation [67], object detection [38] and 3D reconstruction [114]. Finally, the usage of FSL has also been extended to video inputs [234].
5.2 Reinforcement Learning
@@
One of the first applications for many of the aforementioned meta-learning techniques has been meta-RL, which involves learning to quickly adapt across various similar environments by finding an optimal policy across tasks, given only a few episodes in each one of them. This is a particularly useful skill for agents acting in dynamic environments, largely resembling human-like learning, and has been used in the context of imitation learning [170], visual navigation [152], multi-armed bandits [194], robot arm manipulation [212] and control [152]. Novel applications of meta-RL also include online meta-RL [102], black-box optimization [137], long-term memory from visual observations [14] and implicit language learning [131].
5.3 Natural Language Processing
@@
Another field of research where FSL has drawn increasing attention in the last few years is that of NLP, with some standard applications including machine translation [62], and text [73] and sentiment [242] classification. More recently, FSL has also been extended to novel applications and tasks such as prompt tuning [93], visual question answering [149], and dense retrieval [35]. Finally, based on LLMs’ ability to handle multiple downstream tasks when given only a few examples in the form of prompts (see Section 3.4), various similar FSL techniques have also been proposed for applications such as natural language inference including multiple downstream tasks like causal reasoning, textual entailment, word sense disambiguation, and question answering [148].
5.4 Audio